Problem 29.32 when a car drives vertical G42 Part
Answer #1

Since antenna is given it is vertical, the maximum emf will occur when car is traveling perpendicular to the horizontal component of the Earth magnetic field.This occurs when the car is traveling in the east or west direction

Emf = Blv cos (45°)

Given B = 5*10^-4 , l= 64.2 cm ,v = 40.3m/s

= 5*10^-4*0.642 * 40.3 * cos45°

= 91.47*10^-4 V (or ) 9.147 mV

Part A ) emf = 9.147 mV

Part B) option B) car direction should be in east or west